A random sample of 8 applicants for a certain foreign job is selected. The number of years that these applicants have studied French in high school or university, X, and the mark which they obtained in a proficiency test in French, Y, are recorded. The following information is given: ∑X =32 ∑Y =528 ∑XY=2136 ∑X2 =130 ∑Y2 =34872 Fitting a regression line, which of the following statements is incorrect?


1
Expert's answer
2021-08-11T12:39:26-0400

"b=\frac{8*2136-32*528}{8*130-32^2}=\frac{192}{16}=12."

"a=\frac{528-12*32}{8}=\frac{144}{8}=18."

Regression line: "y=a+bx=18+12x."